# On estimation and order selection for multivariate extremes via clustering

### v1

@@ 3.1 The method
Motivated by these observations, we propose to introduce a penalty term that discourages small cluster size and small $D$ dissimilarity between nearest centers. There is arguably some arbitrariness in the choice of this penalty. Through some mathematical heuristics and extensive experiments, we find that the following penalty works relatively well. Let $t\geq 0$ be a tuning parameter. Set
@@
where $\min_{1\leq i<j\leq k}D\left(\mathbf{a}_{i}^{*},\mathbf{a}_{j}^{*}\right)$ is understood as $1$ when $k=1$. Note that $\min_{1\leq i\leq k}|C_{i}|\leq|W|/k$, which explains the normalization in the first denominator above. Recall also that $D$ maps surjectively onto $[0,1]$. Then we form the penalized ASW defined by
@@
Suppose that the assumption in Corollary 2.5 holds, where the true order is $k\in\mathbb{Z}_{+}$ in the discrete spectral measure $H=\sum_{i=1}^{k}p_{i}\delta_{\mathbf{a}_{i}}$ in (18). Let $r_{A}$ be defined as in (20) and define
@@
Suppose $(A_{m,n},\mathfrak{C}_{m,n})$ is an $m$-clustering of $W_{n}$, $m\in\mathbb{Z}_{+}$. Then for any $t\in(0,t_{0})$, where $t_{0}:=\ln\left(1-r_{A}p_{\min}\right)/\ln\left(r_{A}kp_{\min}\right)$, we have
@@
In practice, we suggest plotting the penalized ASW $S_{t}$ as a function of $m=1,2,\ldots$, for a range of small $t$ values starting from $t=0$. We increase $t$ until the curves start to show an obvious upward bend. We then identify the turning point $m$ as the choice of the order $k$. As a quick illustration, we follow a simulation setup of $(d=6,k=6)$ described in 6.1 below to simulate a max-linear factor model (Section 5.1). See Figure 1. It would be desirable to develop a fully automatic data-driven method for choosing $t$, which we leave for a future work to explore.
Fig. 1: A simulation instance taken from Section 6.1 $d=6$, $k=6$ setup. Penalized Average Silhouette Width (ASW) $S_{t}$ (vertical axis) for spherical $k$-means clustering is plotted as a function of test order $m$ (horizontal axis). The different penalty values of $t$ are illustrated by different colors. The true discrete spectral measure in (2.5) is given by $(\mathbf{a}_{1},p_{1})=((0.29,0.21,0.50,0.45,0.43,0.49)^{\top},0.22)$, $(\mathbf{a}_{2},p_{2})=((0.74,0.00,0.59,0.00,0.32,0.00)^{\top},0.10)$, $(\mathbf{a}_{3},p_{3})=((0.00,0.27,0.00,0.47,0.00,0.84)^{\top},0.13)$, $(\mathbf{a}_{4},p_{4})=((0.33,0.70,0.63,0.00,0.00,0.00)^{\top},0.14)$, $(\mathbf{a}_{5},p_{5})=((0.00,0.00,0.00,0.81,0.47,0.34)^{\top},0.09)$, $(\mathbf{a}_{6},p_{6})=((0.48,0.49,0.25,0.33,0.53,0.29)^{\top},0.32)$.
@@ 3.2.1 Some deterministic estimates
We first prepare some deterministic estimates regarding the $k$-clustering in Definition 2.2 and ASW. We shall need the setup in Assumption 2 below in the subsequent developments.
Assumption 2.
Suppose $D$ is a dissimilarity measure that satisfies Assumption 1 and $W$ is a multiset on $\mathbb{S}_{+}^{d-1}$ with $|W|<\infty$. Let $A=\left\{\mathbf{a}_{1},\ldots,\mathbf{a}_{k}\right\}$ be a set of $k$ distinct points on $\mathbb{S}_{+}^{d-1}$ and $p_{i}>0$ with $p_{1}+\ldots+p_{k}=1$, $k\in\mathbb{Z}_{+}$, $k\leq|W|$. Let $\epsilon\in(0,r_{A})$ with $r_{A}$ in (20), and $\delta\in\left(0,p_{\min}\right)$, with $p_{\min}$ in (24), and suppose
$\frac{|W\cap B(\mathbf{a}_{i},\epsilon)|}{|W|}\geq p_{i}-\delta,\quad i\in\left\{1,\ldots,k\right\}.$ (25)
Set for $s>0$ that
$r_{A}^{\dagger}(s)=\sup\left\{D^{\dagger}(\mathbf{a}_{i},\mathbf{w}):i\in\left\{1,\ldots,k\right\},\ \mathbf{w}\in B(\mathbf{a}_{i},s)\right\}.$ (26)
Note that $r_{A}^{\dagger}(s)>0$ for any $s>0$, and $r_{A}^{\dagger}(s)\rightarrow 0$ as $s\rightarrow 0$; see Remark 2.1. The next lemma provides an upper bound for the ASW $\bar{S}$ in (21) when the number of clusters is less than $k$ in Assumption 2.
@@
Suppose Assumptions 1 and 2 hold and $1\leq m<k$. Let $(A^{*}_{m},\mathfrak{C}_{m})$ be an $m$-clustering of $W$. Assume in addition $r_{A}^{\dagger}(\epsilon)<r_{A}$. Then the (unpenalized) ASW $\bar{S}$ satisfies
$\bar{S}=\bar{S}(W;A^{*}_{m},\mathfrak{C}_{m})\leq 1-\left(p_{\min}-\delta\right)\left(r_{A}-r_{A}^{\dagger}(\epsilon)\right).$
@@
Since $m<k$ and $B(\mathbf{a}_{i},r_{A})$’s are disjoint, $i\in\left\{1,\ldots,k\right\}$, there exists $\ell\in\{1,\ldots,k\}$ such that $B(\mathbf{a}_{\ell},r_{A})\cap A_{m}^{*}=\emptyset$ Hence for any $\mathbf{w}\in W\cap B(\mathbf{a}_{\ell},\epsilon)$, we have by the triangular inequality (10) that
@@
$\frac{1}{|W|}\sum_{\mathbf{w}\in W}\frac{a(\mathbf{w})}{b(\mathbf{w})}\geq\frac{1}{|W|}\sum_{\mathbf{w}\in W}a(\mathbf{w})\geq\frac{|W\cap B(\mathbf{a}_{\ell},\epsilon)|}{|W|}\left(r_{A}-r_{A}^{\dagger}(\epsilon)\right)\geq\left(p_{\min}-\delta\right)\left(r_{A}-r_{A}^{\dagger}(\epsilon)\right),$
@@
The next lemma states that when the number of clusters is at least $k$, there will exist at least $k$ centers which are close to $A$ in Assumption 2.
@@
Suppose Assumptions 1 and 2 hold. Let $\left(A^{*}_{m}=\{\mathbf{a}_{1}^{*},\ldots,\mathbf{a}_{m}^{*}\},\mathfrak{C}_{m}\right)$ be an $m$-clustering of $W$, $m\geq k$. Then for any $i\in\{1,\ldots,k\}$, there exists $j\in\{1,\ldots,m\}$, such that $D\left(\mathbf{a}_{j}^{*},\mathbf{a}_{i}\right)<\epsilon^{\prime}$, where
@@
In particular, when $m=k$, and $\epsilon$ and $\delta$ are small enough so that $\epsilon^{\prime}<r_{A}$, there exists a bijection $\pi:\{1,\ldots,k\}\mapsto\{1,\ldots,k\}$, such that $D\left(\mathbf{a}_{\pi(i)}^{*},\mathbf{a}_{i}\right)<\epsilon^{\prime}$ for all $i\in\{1,\ldots,k\}$.
@@
We prove by contradiction. Suppose there exists $i\in\{1,\ldots,k\}$ such that $D\left(\mathbf{a}_{j}^{*},\mathbf{a}_{i}\right)\geq\epsilon^{\prime}$ for all $j\in\{1,\ldots,m\}$. Then for any $\mathbf{w}\in W\cap B(\mathbf{a}_{i},\epsilon)$, we have by the triangular inequality (10) that
@@
Hence combining this and Assumption 2,
$\frac{1}{|W|}\sum_{\mathbf{w}\in W}D\left(\mathbf{w},A_{m}^{*}\right)\geq\frac{1}{|W|}\sum_{\mathbf{w}\in W\cap B(\mathbf{a}_{i},\epsilon)}D\left(\mathbf{w},A_{m}^{*}\right)\geq\left(p_{i}-\delta\right)\left(\epsilon^{\prime}-r_{A}^{\dagger}(\epsilon)\right)\geq\left(p_{\min}-\delta\right)\left(\epsilon^{\prime}-r_{A}^{\dagger}(\epsilon)\right).$ (28)
Next, suppose that a multiset $S$ on $\mathbb{S}_{+}^{d-1}$ contains $A$ and $|S|=m$, which is only possible when $m\geq k$ as assumed. Then we have $D(\mathbf{w},S)\leq D(\mathbf{w},A)$. Set $U_{\epsilon}:=W\cap\left(\cup_{i=1}^{k}B(\mathbf{a}_{i},\epsilon)\right)$, we have that
@@
which contradicts the choice of $\epsilon^{\prime}$. ∎
As a consequence of the previous lemma, when the number of clusters exceeds $k$, either some cluster has a small size or at least two centers are close to each other, as articulated in the next lemma.
@@
Suppose Assumptions 1 and 2 hold. Assume additionally that $\delta$ and $\epsilon$ in Assumption 2 are small enough so that $\epsilon^{\prime}$ in (27) satisfies $\epsilon^{\prime}<r_{A}$. Let $\left(A^{*}_{m}=\{\mathbf{a}_{1}^{*},\ldots,\mathbf{a}_{m}^{*}\},\mathfrak{C}_{m}=\{C_{1},\ldots,C_{m}\}\right)$ be an $m$-clustering of $W$, $m>k$. Then either of the following happens:
@@
Since $B\left(\mathbf{a}_{i},\epsilon^{\prime}\right)$, $i\in\left\{1,\ldots,k\right\}$, are disjoint (because $\epsilon^{\prime}<r_{A}$), by Lemma 3.3, we can, without loss of generality, assume that $\mathbf{a}_{i}^{*}\in B\left(\mathbf{a}_{i},\epsilon^{\prime}\right)$, $i\in\left\{1,\ldots,k\right\}$ We now divide into two cases as follows.
Case 1: there exists one $j\in\{k+1,\ldots,m\}$ (fixed below in the discussion of this case) which satisfies $D\left(\mathbf{a}_{j}^{*},A\right)>\epsilon^{\prime}+2r_{A}^{\dagger}(\epsilon)$. Then for any $i\in\{1,\ldots,k\}$ and any $\mathbf{w}\in W\cap B\left(\mathbf{a}_{i},\epsilon\right)$, we have by the triangular inequality (10) that $D\left(\mathbf{w},\mathbf{a}_{i}^{*}\right)\leq D\left(\mathbf{a}_{i},\mathbf{a}_{i}^{*}\right)+D^{\dagger}\left(\mathbf{a}_{i},\mathbf{w}\right)\leq\epsilon^{\prime}+r_{A}^{\dagger}(\epsilon)$, and hence
@@
This in view of Definition 2.2 implies that $W\cap B\left(\mathbf{a}_{i},\epsilon\right)\subset W\cap C_{j}^{c}$ for all $i\in\left\{1,\ldots,k\right\}$. Therefore, we have by Assumption 2 that
$\min_{i=1,\ldots,m}\frac{|C_{i}|}{|W|}\leq\frac{|C_{j}|}{|W|}\leq\frac{|W\cap\bigcap_{i=1,\ldots,k}B\left(\mathbf{a}_{i},\epsilon\right)^{c}|}{|W|}\leq k\delta.$
Case 2: for any $j\in\{k+1,\ldots,m\}$, we have $D\left(\mathbf{a}_{j}^{*},\mathbf{a}_{i}\right)\leq\epsilon^{\prime}+2r_{A}^{\dagger}(\epsilon)$ for all $i\in\{1,\ldots,k\}$. Then for any such pair of $j$ and $i$, we have
@@
In the scenario where the number of clusters matches the specified value $k$ in Assumption 2, the next lemma establishes a lower bound for the (unpenalized) ASW $\bar{S}$. Furthermore, it provides lower bounds for both the sizes of individual clusters and the dissimilarities between cluster centers.
@@
Suppose Assumptions 1 and 2 hold. Let $\left(A^{*}_{k}=\{\mathbf{a}_{1}^{*},\ldots,\mathbf{a}_{k}^{*}\},\mathfrak{C}_{k}=\{C_{1},\ldots,C_{k}\}\right)$ be a $k$-clustering of $W$. Suppose in addition
@@
$\bar{S}=\bar{S}(W_{n};A^{*}_{k},\mathfrak{C}_{k})\geq 1-\left(1-k\delta\right)\frac{\epsilon^{\prime}+r_{A}^{\dagger}(\epsilon)}{r_{A}-r_{A}^{\dagger}(\epsilon)-r_{A}^{\dagger}(\epsilon^{\prime})}-k\delta.$
@@
Since $r_{A}>\epsilon^{\prime}$, by Lemma 3.3, there exists a permutation $\pi:\{1,\ldots,k\}\mapsto\{1,\ldots,k\}$, such that $D\left(\mathbf{a}_{i},\mathbf{a}_{\pi(i)}^{*}\right)<\epsilon^{\prime}$, $i\in\left\{1,\ldots,k\right\}$. Then for each $i$ and any $\mathbf{w}\in B(\mathbf{a}_{i},\epsilon)$, we have by the triangular inequality (10) that
@@
where if $k=1$, the left-hand side $D\left(\mathbf{w},\mathbf{a}_{\pi(j)}^{*}\right)$ in (32) is understood as 1, and the inequality still holds. Writing as before $U_{\epsilon}=\bigcup_{1\leq i\leq k}B(\mathbf{a}_{i},\epsilon)\cap W$. In view of Assumption 2 and the inequalities above, we have
@@
For the second claim, in view of Definition 2.2, (30), (31), (32), we have $W\cap B\left(\mathbf{a}_{i},\epsilon\right)\subset C_{\pi(i)}$, $i\in\left\{1,\ldots,k\right\}$. Hence by Assumption 2,
$\frac{|C_{\pi(i)}|}{|W|}\geq\frac{|W\cap B\left(\mathbf{a}_{i},\epsilon\right)|}{|W|}\geq p_{i}-\delta.$
@@
Following the setup and notation of Sections 2.1 and Section 2.3, recall $W_{n}$ denotes the extremal subsample as in (14), and ${A}_{m,n}=\left\{{\mathbf{a}}_{1,n}^{m},\ldots,{\mathbf{a}}_{m,n}^{m}\right\}$, ${C}_{i,n}^{m}$, $i\in\{1,\ldots,m\}$, are random multisets on $\mathbb{S}_{+}^{d-1}$ such that $\left(A_{m,n},\mathfrak{C}_{m,n}=\{{C}_{1,n}^{m},\ldots,{C}_{m,n}^{m}\}\right)$ form an $m$-clustering of $W_{n}$, $m\in\mathbb{Z}_{+}$, $n\geq m$. Throughout this section, we follow the assumption of a discrete spectral measure as in Corollary 2.5, and suppose the dissimilarity measure $D$ satisfies Assumption 1.
@@
If $m<k$, then almost surely,
$\limsup_{n}\bar{S}\left(W_{n};A_{m,n},\mathfrak{C}_{m,n}\right)\leq 1-r_{A}p_{\min}.$
If $m=k$, then almost surely,

### v2

@@ 3.1 The method
Motivated by these observations, we propose to introduce a penalty term that discourages small cluster size and small $D$ dissimilarity between centers. Recall that for a $k$-clustering $\left(A_{k}^{*}=\left\{\mathbf{a}_{1}^{*},\ldots,\mathbf{a}_{k}^{*}\right\},\mathfrak{C}_{k}=\left\{C_{1},\ldots,C_{k}\right\}\right)$ of a multiset $W$, the set $\left\{\mathbf{a}_{1}^{*},\ldots,\mathbf{a}_{k}^{*}\right\}$ represents the cluster centers and $\left\{C_{1},\ldots,C_{k}\right\}$ denotes the partition of $W$. Then $\min_{i=1,\ldots,k}\left(\frac{|C_{i}|}{|W|/k}\right)$ is the smallest cluster size relative to the average cluster size $|W|/k$, and $\min_{1\leq i<j\leq k}D\left(\mathbf{a}_{i}^{*},\mathbf{a}_{j}^{*}\right)$ is the smallest pairwise dissimilarity between centers. Note that both quantities are between 00 and $1$ (recall that $D$ maps onto $[0,1]$), and we want to introduce a criterion that penalizes either of them being small. There is arguably some arbitrariness in designing this penalty criterion. Through some mathematical heuristics and extensive experiments, we find that the following penalty works relatively well. Let $t\geq 0$ be a tuning parameter. Set
@@
where $\min_{1\leq i<j\leq k}D\left(\mathbf{a}_{i}^{*},\mathbf{a}_{j}^{*}\right)$ is understood as $1$ when $k=1$. Then we form the penalized ASW defined by
@@
Suppose $\mathbf{X}$ satisfying (1) and (2) has a discrete spectral measure of the form $H=\sum_{i=1}^{k}p_{i}\delta_{\mathbf{a}_{i}}$, $k\in\mathbb{Z}_{+}$ where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\ldots+p_{k}=1$. Let $W_{n}$ denote the extremal subsample as in (14) with $\ell_{n}\rightarrow\infty$, $\ell_{n}/n\rightarrow 0$, and $\left(A_{m,n},\mathfrak{C}_{m,n}\right)$, $m\in\mathbb{Z}_{+}$, form an $m$-clustering of $W_{n}$ as defined in Definition 2.3 with respect to a dissimilarity measure $D$ defined in Definition 2.1. Let $r_{A}$ be defined as in (20) and define
@@
Then for any $t\in(0,t_{0})$, where
$t_{0}:=\ln\left(1-r_{A}p_{\min}\right)/\ln\left(r_{A}kp_{\min}\right),$ (25)
we have
@@
In practice, we suggest plotting the penalized ASW $S_{t}$ as a function of $m=1,2,\ldots$, for a range of small $t$ values. The idea is to start with $t$ near 00, gradually increase it, and see how the penalized ASW curve responds. If some obvious spurious clusters (i.e., those with tiny size or centers that are too close together) are present, the curve tends to respond sensitively and bends at the appropriate order. We then identify the turning point $m$ as the choice of the order $k$. As a quick illustration, we follow a simulation setup of $(d=6,k=6)$ described in Section 6.1 below to simulate a max-linear factor model (Section 5.1). See Figure 1. Increasing $t$ to a very large value is not informative and is not recommended in practice. On the other hand, it would be desirable to develop a data-driven method for choosing $t$, which we leave for a future work to explore.
Fig. 1: A simulation instance taken from Section 6.1 $d=6$, $k=6$ setup. Penalized Average Silhouette Width (ASW) $S_{t}$ (vertical axis) for spherical $k$-means clustering is plotted as a function of test order $m$ (horizontal axis). The different penalty values of $t$ are illustrated by different colors. The true discrete spectral measure in (18) is given by $(\mathbf{a}_{1},p_{1})=((0.29,0.21,0.50,0.45,0.43,0.49)^{\top},0.22)$, $(\mathbf{a}_{2},p_{2})=((0.74,0.00,0.59,0.00,0.32,0.00)^{\top},0.10)$, $(\mathbf{a}_{3},p_{3})=((0.00,0.27,0.00,0.47,0.00,0.84)^{\top},0.13)$, $(\mathbf{a}_{4},p_{4})=((0.33,0.70,0.63,0.00,0.00,0.00)^{\top},0.14)$, $(\mathbf{a}_{5},p_{5})=((0.00,0.00,0.00,0.81,0.47,0.34)^{\top},0.09)$, $(\mathbf{a}_{6},p_{6})=((0.48,0.49,0.25,0.33,0.53,0.29)^{\top},0.32)$.
@@ 3.2.1 Some deterministic estimates
We first prepare some deterministic estimates regarding the $k$-clustering in Definition 2.3 and ASW in (21). Based on Proposition 2.5, if the true spectral measure $H$ consists of finitely many atoms, and we select the subset $S$ in the proposition as a union of neighborhoods around each atom, then, for sufficiently large $n$, almost all points in the extremal subsample $W_{n}$ in (14) will lie near one of these atoms with respect to the dissimilarity measure $D$. This motivates us to address a scenario where we aim to cluster a multiset where the majority of points are concentrated around a finite number of centers.
Fix throughout this section distinct points $\left\{\mathbf{a}_{1},\ldots,\mathbf{a}_{k}\right\}=:A$ on $\mathbb{S}_{+}^{d-1}$, $k\in\mathbb{Z}_{+}$, $p_{i}>0$ with $p_{1}+\ldots+p_{k}=1$, and a dissimilarity measure $D$ defined in Definition 2.1. We say a multiset $W$ on $\mathbb{S}_{+}^{d-1}$ with $|W|<\infty$ satisfies the concentration condition $\mathcal{A}(\epsilon,\delta)$ with $\epsilon,\delta>0$, if $k\leq|W|\wedge\delta^{-1}$ such that $\epsilon\in(0,r_{A})$ with $r_{A}$ in (20), $\delta\in\left(0,p_{\min}\right)$ with $p_{\min}$ in (24), and
$\frac{|W\cap B_{D}(\mathbf{a}_{i},\epsilon)|}{|W|}\geq p_{i}-\delta,\quad i\in\left\{1,\ldots,k\right\}.$ (26)
In proving the following lemmas, we need to control the dual dissimilarity $D^{\dagger}$ of points within the same $D$-neighborhood. Therefore, we introduce the following uniform upper bound. Set for the aforementioned $A=\left\{\mathbf{a}_{1},\ldots,\mathbf{a}_{k}\right\}$ and $s>0$ that
$r_{A}^{\dagger}(s)=\sup\left\{D^{\dagger}(\mathbf{a}_{i},\mathbf{w}):i\in\left\{1,\ldots,k\right\},\ \mathbf{w}\in B_{D}(\mathbf{a}_{i},s)\right\}.$ (27)
Remark 3.2.
Note that $r_{A}^{\dagger}(s)>0$ for any $s>0$, and $r_{A}^{\dagger}(s)\rightarrow 0$ as $s\rightarrow 0$; see Remark 2.2. For sufficiently small $\epsilon$, we always have $r_{A}^{\dagger}(\epsilon)<r_{A}$.
If $W$ satisfies the concentration condition $\mathcal{A}(\epsilon,\delta)$ but is partitioned into fewer clusters than $k$, then at least two of the true clusters will be merged into a single cluster. This merge can be effectively identified by the ASW based on the reduced separability, as demonstrated in the next lemma.
@@
Suppose a multiset $W$ satisfies the condition $\mathcal{A}(\epsilon,\delta)$ and $1\leq m<k$. Let $(A^{*}_{m},\mathfrak{C}_{m})$ be an $m$-clustering of $W$ as defined in Definition 2.3. Then the (unpenalized) ASW $\bar{S}$ satisfies
$\bar{S}=\bar{S}(W;A^{*}_{m},\mathfrak{C}_{m})\leq 1-\left(p_{\min}-\delta\right)\left(r_{A}-r_{A}^{\dagger}(\epsilon)\right),$
where $r_{A}^{\dagger}$ is as in (27).
@@
Since $m<k$ and $B_{D}(\mathbf{a}_{i},r_{A})$’s are disjoint, $i\in\left\{1,\ldots,k\right\}$, there exists $\ell\in\{1,\ldots,k\}$ such that $B_{D}(\mathbf{a}_{\ell},r_{A})\cap A_{m}^{*}=\emptyset$. Hence for any $\mathbf{w}\in W\cap B_{D}(\mathbf{a}_{\ell},\epsilon)$, we have by the triangular inequality (10) that
@@
$\frac{1}{|W|}\sum_{\mathbf{w}\in W}\frac{a(\mathbf{w})}{b(\mathbf{w})}\geq\frac{1}{|W|}\sum_{\mathbf{w}\in W}a(\mathbf{w})\geq\frac{|W\cap B_{D}(\mathbf{a}_{\ell},\epsilon)|}{|W|}\left(r_{A}-r_{A}^{\dagger}(\epsilon)\right)\geq\left(p_{\min}-\delta\right)\left(r_{A}-r_{A}^{\dagger}(\epsilon)\right),$
@@
The next lemma states that if $W$ satisfies the concentration condition $\mathcal{A}(\epsilon,\delta)$ and is partitioned into a number of clusters that is greater than or equal to $k$, then for each true center of $W$, there exists at least one cluster center that is close to it with respect to the dissimilarity measure $D$.
@@
Suppose a multiset $W$ satisfies the condition $\mathcal{A}(\epsilon,\delta)$. Let $\left(A^{*}_{m}=\{\mathbf{a}_{1}^{*},\ldots,\mathbf{a}_{m}^{*}\},\mathfrak{C}_{m}\right)$ be an $m$-clustering of $W$ as defined in Definition 2.3, $m\geq k$. Then for any $i\in\{1,\ldots,k\}$, there exists $j\in\{1,\ldots,m\}$, such that $D\left(\mathbf{a}_{j}^{*},\mathbf{a}_{i}\right)<\epsilon^{\prime}$, where
@@
In particular, when $m=k$ and $\epsilon^{\prime}<r_{A}$, there exists a bijection $\pi:\{1,\ldots,k\}\mapsto\{1,\ldots,k\}$, such that $D\left(\mathbf{a}_{\pi(i)}^{*},\mathbf{a}_{i}\right)<\epsilon^{\prime}$ for all $i\in\{1,\ldots,k\}$.
@@
We prove the first claim by contradiction. Suppose there exists $i\in\{1,\ldots,k\}$ such that $D\left(\mathbf{a}_{j}^{*},\mathbf{a}_{i}\right)\geq\epsilon^{\prime}$ for all $j\in\{1,\ldots,m\}$. Then for any $\mathbf{w}\in W\cap B_{D}(\mathbf{a}_{i},\epsilon)$, we have by the triangular inequality (10) that
@@
Hence combining this and (26),
$\frac{1}{|W|}\sum_{\mathbf{w}\in W}D\left(\mathbf{w},A_{m}^{*}\right)\geq\frac{1}{|W|}\sum_{\mathbf{w}\in W\cap B_{D}(\mathbf{a}_{i},\epsilon)}D\left(\mathbf{w},A_{m}^{*}\right)\geq\left(p_{i}-\delta\right)\left(\epsilon^{\prime}-r_{A}^{\dagger}(\epsilon)\right)\geq\left(p_{\min}-\delta\right)\left(\epsilon^{\prime}-r_{A}^{\dagger}(\epsilon)\right).$ (29)
Next, suppose that a multiset $S$ on $\mathbb{S}_{+}^{d-1}$ contains $A$ and $|S|=m$, which is only possible when $m\geq k$ as assumed. Then we have $D(\mathbf{w},S)\leq D(\mathbf{w},A)$. Set $U_{\epsilon}:=W\cap\left(\cup_{i=1}^{k}B_{D}(\mathbf{a}_{i},\epsilon)\right)$, we have that
@@
which contradicts the choice of $\epsilon^{\prime}$.
For the second claim, note that $B_{D}(\mathbf{a}_{i},r_{A})$’s are disjoint, $i\in\left\{1,\ldots,k\right\}$. So if $\epsilon^{\prime}<r_{A}$, it is impossible that $D\left(\mathbf{a}_{j}^{*},\mathbf{a}_{i}\right)<\epsilon^{\prime}$ and $D\left(\mathbf{a}_{j}^{*},\mathbf{a}_{i^{\prime}}\right)<\epsilon^{\prime}$ hold simultaneously when $i\neq i^{\prime}$. The conclusion then follows. ∎
As a consequence of the previous lemma, if $W$ is concentrated around $k$ centers but is partitioned into more than $k$ clusters, then either some clusters will have a small size, or at least two of the centers will be close to each other, as articulated in the next lemma.
@@
Suppose a multiset $W$ satisfies the condition $\mathcal{A}(\epsilon,\delta)$. Assume additionally that $\epsilon^{\prime}$ in (28) satisfies $\epsilon^{\prime}<r_{A}$. Let $\left(A^{*}_{m}=\{\mathbf{a}_{1}^{*},\ldots,\mathbf{a}_{m}^{*}\},\mathfrak{C}_{m}=\{C_{1},\ldots,C_{m}\}\right)$ be an $m$-clustering of $W$ as defined in Definition 2.3, $m>k$. Then either of the following happens:
@@
Since $B_{D}\left(\mathbf{a}_{i},\epsilon^{\prime}\right)$, $i\in\left\{1,\ldots,k\right\}$, are disjoint (because $\epsilon^{\prime}<r_{A}$), by Lemma 3.4, we can, without loss of generality, assume that $\mathbf{a}_{i}^{*}\in B_{D}\left(\mathbf{a}_{i},\epsilon^{\prime}\right)$, $i\in\left\{1,\ldots,k\right\}$ We now divide into two cases as follows.
Case 1: there exists one $j\in\{k+1,\ldots,m\}$ (fixed below in the discussion of this case) which satisfies $D\left(\mathbf{a}_{j}^{*},A\right)>\epsilon^{\prime}+2r_{A}^{\dagger}(\epsilon)$. Then for any $i\in\{1,\ldots,k\}$ and any $\mathbf{w}\in W\cap B_{D}\left(\mathbf{a}_{i},\epsilon\right)$, we have by the triangular inequality (10) that $D\left(\mathbf{w},\mathbf{a}_{i}^{*}\right)\leq D\left(\mathbf{a}_{i},\mathbf{a}_{i}^{*}\right)+D^{\dagger}\left(\mathbf{a}_{i},\mathbf{w}\right)\leq\epsilon^{\prime}+r_{A}^{\dagger}(\epsilon)$, and hence
@@
This in view of Definition 2.3 implies that $W\cap B_{D}\left(\mathbf{a}_{i},\epsilon\right)\subset W\cap C_{j}^{c}$ for all $i\in\left\{1,\ldots,k\right\}$. Therefore, we have by (26) that
$\min_{i=1,\ldots,m}\frac{|C_{i}|}{|W|}\leq\frac{|C_{j}|}{|W|}\leq\frac{|W\cap\bigcap_{i=1,\ldots,k}B_{D}\left(\mathbf{a}_{i},\epsilon\right)^{c}|}{|W|}\leq k\delta.$
Case 2: for any $j\in\{k+1,\ldots,m\}$, we have $D\left(\mathbf{a}_{j}^{*},\mathbf{a}_{i}\right)\leq\epsilon^{\prime}+2r_{A}^{\dagger}(\epsilon)$ for some $i\in\{1,\ldots,k\}$. Then for any such pair of $j$ and $i$, we have
@@
The next lemma says, if $W$ is concentrated around $k$ centers and is partitioned into exactly $k$ clusters, the (unpenalized) ASW $\bar{S}$ will favor this clustering by providing a high score that is close to $1$. Additionally, none of the clusters will have a size too small, and no two cluster centers will be too close to each other.
@@
Suppose a multiset $W$ satisfies the condition $\mathcal{A}(\epsilon,\delta)$. Let $\left(A^{*}_{k}=\{\mathbf{a}_{1}^{*},\ldots,\mathbf{a}_{k}^{*}\},\mathfrak{C}_{k}=\{C_{1},\ldots,C_{k}\}\right)$ be a $k$-clustering of $W$ as defined in Definition 2.3. Suppose in addition
@@
$\bar{S}=\bar{S}(W;A^{*}_{k},\mathfrak{C}_{k})\geq 1-\left(1-k\delta\right)\frac{\epsilon^{\prime}+r_{A}^{\dagger}(\epsilon)}{r_{A}-r_{A}^{\dagger}(\epsilon)-r_{A}^{\dagger}(\epsilon^{\prime})}-k\delta.$
@@
Since $r_{A}>\epsilon^{\prime}$, by Lemma 3.4, there exists a permutation $\pi:\{1,\ldots,k\}\mapsto\{1,\ldots,k\}$, such that $D\left(\mathbf{a}_{i},\mathbf{a}_{\pi(i)}^{*}\right)<\epsilon^{\prime}$, $i\in\left\{1,\ldots,k\right\}$. Then for each $i$ and any $\mathbf{w}\in B_{D}(\mathbf{a}_{i},\epsilon)$, we have by the triangular inequality (10) that
@@
where if $k=1$, the left-hand side $D\left(\mathbf{w},\mathbf{a}_{\pi(j)}^{*}\right)$ in (33) is understood as 1, and the inequality still holds. Writing as before $U_{\epsilon}=\bigcup_{1\leq i\leq k}B_{D}(\mathbf{a}_{i},\epsilon)\cap W$. In view of (26) and the inequalities above, we have
@@
For the second claim, in view of Definition 2.3, (31), (32), (33), we have $W\cap B_{D}\left(\mathbf{a}_{i},\epsilon\right)\subset C_{\pi(i)}$, $i\in\left\{1,\ldots,k\right\}$. Hence by (26),
$\frac{|C_{\pi(i)}|}{|W|}\geq\frac{|W\cap B_{D}\left(\mathbf{a}_{i},\epsilon\right)|}{|W|}\geq p_{i}-\delta.$
@@
Suppose $\mathbf{X}$ satisfying (1) and (2) has a discrete spectral measure of the form $H=\sum_{i=1}^{k}p_{i}\delta_{\mathbf{a}_{i}}$, where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\ldots+p_{k}=1$. Let $W_{n}$ denote the extremal subsample as in (14), and $\left(A_{m,n},\mathfrak{C}_{m,n}\right)$ form an $m$-clustering of $W_{n}$ as defined in Definition 2.3 with respect to a dissimilarity measure $D$ defined in Definition 2.1. If $m<k$, then almost surely,
$\limsup_{n}\bar{S}\left(W_{n};A_{m,n},\mathfrak{C}_{m,n}\right)\leq 1-r_{A}p_{\min},$
where $r_{A}$ is as in (20) and $p_{\min}$ is as in (24). If $m=k$, then almost surely,